# Low-Rank Bandits via Tight Two-to-Infinity Singular Subspace Recovery

### v1

@@ C.3 Additional Lemmas
for some universal constant $c>0$. Recall that we denoted by $A_{\delta}$ the event under which the bounds of Lemma D.5 holds: under $A_{\delta}$,
$\epsilon$ $\leq\frac{C_{0}\|M\|_{\max}\mu^{6}\kappa^{4}r^{3}\left(m+n\right)\log^{3}\left((m+n)T_{1}/\delta\right)}{T_{1}\omega_{\min}\sqrt{mn}\left(m\wedge n\right)},$
$L_{\theta}$ $\leq C_{1}\|M\|^{2}_{\max}\mu^{4}\kappa^{4}r^{2}\frac{m\vee n}{m\wedge n},$
@@
for some universal constants $C_{0},C_{1},C_{2}.$ Note that $\frac{1}{\sqrt{T_{2}}}\leq\frac{1}{\sqrt{\left(1-\alpha\right)T}}$. By Lemma D.3, under $A_{\delta}$, with probability larger than $21-\delta/$ 2, we have
$\left|L_{1}\right|+\left|L_{2}\right|\leq\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}+\operatorname{poly}\left(m,n,\omega_{\min}^{-1},\mu,\kappa,r,\|M\|_{\max}\right)\frac{\log^{2}\left(e/\delta\right)}{T},$
$|L_{3}|\leq\operatorname{poly}\left(m,n,\omega_{\min}^{-1},\mu,\kappa,r,\|M\|_{\max}\right)\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}.$
A union bound concludes the proof of the first part of the proposition.
Let us now upper bound the polynomial term. Direct computations ensure that when the high probability bounds of Lemma D.3 and D.5 hold,
$\left|L_{3}\right|\lesssim\frac{\|M\|_{\max}\mu^{8}\kappa^{5}r^{4}\left(m+n\right)^{2}\log^{3}\left(\left(m+n\right)T/\delta\right)}{T\left(mn\right)^{3/2}\omega_{\min}^{2}\left(m\wedge n\right)},$
$\frac{E_{2}}{T}\lesssim\frac{\max\left(1,\|M\|_{\max}\right)^{2}\mu^{7}\kappa^{11/2}r^{7/2}\left(m+n\right)^{3/2}\left(m\vee n\right)\log\left(ed/\delta\right)}{T\left(\omega_{\min}mn\right)^{3/2}\left(m\wedge n\right)},$
$\frac{E_{3}}{T^{3/2}}$ $\lesssim\frac{\max\left(1,\|M\|_{\max}\right)^{2}\mu^{8}\kappa^{6}r^{4}\left(m+n\right)^{2}\left(m\vee n\right)\log^{3/2}\left(ed/\delta\right)}{T^{3/2}\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}$
$\lesssim\frac{\max\left(1,\|M\|_{\max}\right)^{2}\mu^{5}\kappa^{5}r^{3}\left(m+n\right)^{3/2}\left(m\vee n\right)}{T\left(\omega_{\min}mn\right)^{3/2}\left(m\wedge n\right)},$
[…]
$\lesssim\frac{\left(m+n\right)^{3/2}}{T\omega_{\min}mn\log\left(ed/\delta\right)}$
where we have used $T\gtrsim\frac{\mu^{4}\kappa^{2}r^{2}(m+n)\log^{3}\left(ed/\delta\right)}{\omega_{\min}mn}$ in the last inequalities. Consequently, we can take
$\operatorname{poly}\left(m,n,\omega_{\min}^{-1},\mu,\kappa,r,\|M\|_{\max}\right)=C^{\prime}\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}$
@@
This would yield an error bound whose dominant term scales with $\|\psi_{\pi}\|_{\Lambda}^{-1}\sqrt{\frac{\log\left(1/\delta\right)}{T}}$ and would echo the linear bandit upper bound of Theorem C.7, that we have demonstrated is instance-optimal. Although numerical experiments suggest this result holds, we were unsuccessful in our proof attempts.
By using Lemma D.12, we can however obtain the following instance-dependent bound, with a dependency on $\omega_{\min}$ instead of the entire behavior policy. Theorem 6.2 will follow from this result.
@@
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|$ $\leq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}+J\left(\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}+\frac{\log^{5/4}\left(\left(m+n\right)T/\delta\right)}{T^{3/4}}\right),$
@@
$J=C\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}$ (33)
for some universal constant $C^{\prime}>0$.
@@
$\|\phi_{\pi}\|^{2}_{\Lambda_{\phi}^{-1}}\leq\frac{\|\phi_{\pi}\|^{2}_{2}}{\omega_{\min}}\leq\frac{1}{\omega_{\min}}\left(\|\psi_{\pi}\|^{2}_{2}+3\max\left(\|UU^{\top}-\widehat{U}\widehat{U}^{\top}\|_{2\to\infty},\|VV^{\top}-\widehat{V}\widehat{V}^{\top}\|_{2\to\infty}\right)\right).$
Under the event $A_{\delta}$ defined in the proof of Lemma D.5, we consequently also have
$\|\psi_{\pi}\|^{2}_{\Lambda_{\phi}^{-1}}\leq\frac{1}{\omega_{\min}}\left(\|\psi_{\pi}\|^{2}_{2}+\frac{1}{\sqrt{T}}C\mu^{3}\kappa^{2}r^{3/2}\sqrt{\frac{m+n}{mn\left(m\wedge n\right)\omega_{\min}}}\log^{3/2}\left(\frac{\left(m+n\right)T}{\delta}\right)\right)$ (34)
for some universal constant $C>0$ Since $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ for every $x,y\geq 0$, this implies
$\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\leq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}+C^{1/2}\mu^{3/2}\kappa r^{3/4}\omega_{\min}^{-3/4}\left(\frac{m+n}{mn\left(m\wedge n\right)}\right)^{1/4}\frac{\log^{3/4}\left(\left(m+n\right)T/\delta\right)}{T^{1/4}}.$ (35)
Combining this result with Proposition D.2121212Note that there is no need for a union bound as the high probability bound (35) is true under $A_{\delta}$., we get that with probability larger than $1-\delta,$
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|$ $\leq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}+C_{1}\sqrt{\log(e/\delta)}\mu^{3/2}\kappa r^{3/4}\omega_{\min}^{-3/4}\left(\frac{m+n}{mn\left(m\wedge n\right)}\right)^{1/4}\frac{\log^{3/4}\left(\left(m+n\right)T/\delta\right)}{T^{3/4}}$
$+C_{2}\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$
for some universal constants $C_{1},C_{2}>0$. Since $\sqrt{\log(e/\delta)}\mu^{3/2}\kappa r^{3/4}\omega_{\min}^{-3/4}\left(\frac{m+n}{mn\left(m\wedge n\right)}\right)^{1/4}\log^{3/4}\left(\left(m+n\right)T/\delta\right)\lesssim\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\log^{5/4}\left(\left(m+n\right)T/\delta\right)$, this concludes the proof.
@@
When $M$ and $\omega$ are homogeneous, $\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\lesssim mn$.
When both the context distribution and the behavior policy are uniform, the upper bound of Theorem D.7 corresponds to the conjectured instance-dependent upper bound mentioned earlier in the section. Indeed, Lemma D.11 ensures that in this case, $\Lambda=\frac{1}{mn}I_{d},$ and thus $\|\psi_{\pi}\|_{\Lambda^{-1}}=\sqrt{mn}\|\psi_{\pi}\|_{2}=\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}.$ To derive the tighter instance-dependent bound in the general case, one may be tempted to write $\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}^{2}=\|\psi_{\pi}\|_{\Lambda^{-1}}^{2}+\left(\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}^{2}-\|\phi_{\pi}\|_{\Lambda^{-1}}^{2}\right)+\left(\|\phi_{\pi}\|_{\Lambda^{-1}}^{2}-\|\psi_{\pi}\|_{\Lambda^{-1}}^{2}\right)$ and prove that the two rightmost terms converge to 00 with high probability. However, numerical experiments suggest that these terms do not both vanish in general, but rather cancel out, and thus a more careful analysis is required.
We now discuss how to reformulate the statement of Theorem D.7 as as sample complexity result and retrieve the statement of Theorem 6.2.
@@
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|\leq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}+J\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}+J\frac{\log^{5/4}\left(\left(m+n\right)T/\delta\right)}{T^{3/4}}$
@@
$\frac{\varepsilon}{3}\geq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\log(16/\delta)}{\left(1-\alpha\right)T}}\text{, }\quad\frac{\varepsilon}{3}\geq J\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T},\quad\text{and}\quad\frac{\varepsilon}{3}\geq J\frac{\log^{5/4}\left(\left(m+n\right)T/\delta\right)}{T^{3/4}}.$
@@
$T\gtrsim\frac{\|\psi_{\pi}\|^{2}_{2}}{\omega_{\min}\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{J^{4/3}}{\varepsilon^{4/3}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)\!$
is enough to satisfy all three inequalities, where we recall that $J$ has been defined in (33).
Furthermore, the conditions on the number of samples of Theorem D.7, namely $T\gtrsim\frac{1}{\omega_{\min}mn}\mu^{4}\kappa^{2}r^{2}(m+n)\log^{3}\left((m+n)T/\delta\right)$ and $T\gtrsim\frac{1}{\omega_{\min}}\log\left(16d/\delta\right)$, are both implied by such a choice at the cost of replacing $J^{4/3}$ by $\max\left(1/w_{\min},J^{4/3}\right)$.
@@
$K_{1}:=\max\left(\frac{1}{w_{\min}^{3/4}},\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\right)^{4/3}.$ (36)
We finally note that when $M$ and $\omega$ are homogeneous, $K_{1}=\Theta\left(\left(mn\right)^{4/3}\right).$
Minimax sample complexity guarantee. Let us note that by (41), $\|\psi_{\pi}\|^{2}_{2}\leq\mu^{2}r\frac{m+n-\mu^{2}r}{mn}$, which implies that $\hat{v}_{\text{\tt RS-PE}}$ is $\left(\varepsilon,\delta\right)$-PAC as soon as
$T\gtrsim\mu^{2}r\frac{m+n-\mu^{2}r}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{K_{1}}{\varepsilon^{4/3}}\log^{3}\left(\frac{(m+n)T}{\delta}\right).$
By adapting the previous analysis with a more refined argument, the dependency in $\varepsilon$ and the problem parameters of the second term can be improved. This is what we demonstrate now.
@@
Let $\varepsilon,\delta\in(0,1)$. With the choices $T_{1}=\lfloor\alpha T\rfloor$ and $\tau\leq r(m\wedge n)^{-1}(\omega_{\min}/\omega_{\max})\log(16d/\delta)$, $\hat{v}_{\text{\tt RS-PE}}$ is $\left(\varepsilon,\delta\right)$-PAC as soon as
$T\gtrsim\frac{\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{K_{2}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$
for
$K_{2}:=\max\left(\frac{1}{w_{\min}},\max\left(1,\|M\|_{\max}^{2}\right)\mu^{8}\kappa^{11/2}r^{4}\frac{\left(m+n\right)^{2}\left(m\vee n\right)}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\right)=K_{1}^{3/4}.$ (37)
@@
Note that in the homogeneous case, $K_{2}=\Theta\left(mn\right).$
Proof.
Proof of Theorem D.9 Note that combining (34) with the bound on $\|\psi_{\pi}\|^{2}_{2}$ entails
$\|\psi_{\pi}\|^{2}_{\Lambda_{\phi}^{-1}}$ $\leq\frac{1}{\omega_{\min}}\left(\frac{\mu^{2}r\left(m+n\right)}{mn}+\frac{1}{\sqrt{T}}C\mu^{3}\kappa^{2}r^{3/2}\sqrt{\frac{m+n}{mn\left(m\wedge n\right)\omega_{\min}}}\log^{3/2}\left(\frac{\left(m+n\right)T}{\delta}\right)\right)$
$=\frac{\mu^{2}r\left(m+n\right)}{\omega_{\min}mn}\left(1+\frac{1}{\sqrt{T}}C\mu\kappa^{2}r^{1/2}\sqrt{\frac{mn}{\left(m+n\right)\left(m\wedge n\right)\omega_{\min}}}\log^{3/2}\left(\frac{\left(m+n\right)T}{\delta}\right)\right)$ (38)

### v2

@@ C.3 Additional Lemmas
for some universal constant $c>0$. Recall that we denoted by $A_{\delta}$ the event under which the bounds of Lemma D.5 holds. By Lemma D.3, under $A_{\delta}$, with probability larger than $21-\delta/$ 2, we have
$\left|L_{1}\right|+\left|L_{2}\right|\leq\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\sqrt{\frac{2\sigma^{2}\log(16/\delta)}{\left(1-\alpha\right)T}}+K_{0}\frac{\log^{2}\left(e/\delta\right)}{T},$
$|L_{3}|\leq K_{0}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$
for $K_{0}=\operatorname{poly}\left(\|M\|_{\max},\sigma,\mu,\kappa,r,m,n,(\omega_{\min}mn)^{-1}\right)$ (we have used that $\frac{1}{\sqrt{T_{2}}}\leq\frac{1}{\sqrt{\left(1-\alpha\right)T}}$). A union bound concludes the proof of the first part of Proposition D.2.
Let us now upper bound the polynomial term $K_{0}$. Direct computations ensure that when the high probability bounds of Lemma D.3 and D.5 hold,
$\left|L_{3}\right|$ $\lesssim\epsilon_{\max}L_{\phi}\lesssim\frac{L^{2}}{\sigma_{r}}\frac{\mu^{4}\kappa^{3}r^{2}\left(m+n\right)^{2}\log^{3}\left(\left(m+n\right)T/\delta\right)}{T\omega_{\min}^{2}mn\left(m\wedge n\right)}$
$\lesssim\frac{L^{2}}{\|M\|_{\max}}\frac{\mu^{6}\kappa^{4}r^{3}\left(m+n\right)^{2}\sqrt{mn}\log^{3}\left(\left(m+n\right)T/\delta\right)}{T\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)},$
$\frac{E_{2}}{T}\lesssim\frac{\sqrt{L_{\phi}}\left(L_{\phi}L_{\theta}+\sigma\sqrt{dL_{\phi}}\right)\log(d/\delta)}{T}\lesssim\frac{\left(\|M\|_{\max}+\sigma\right)\mu^{5}r^{5/2}\left(m+n\right)^{3/2}\sqrt{mn}\log\left(d/\delta\right)}{T\left(\omega_{\min}mn\right)^{3/2}\left(m\wedge n\right)},$
$\frac{E_{3}}{T^{3/2}}$ $\lesssim\frac{L_{\phi}\left(\sigma\sqrt{d}+\sqrt{L_{\phi}}L_{\theta}\right)\log(d/\delta)}{T}\lesssim\frac{\left(\|M\|_{\max}+\sigma\right)\mu^{5}r^{5/2}\left(m+n\right)^{3/2}\sqrt{mn}\log\left(d/\delta\right)}{T\left(\omega_{\min}mn\right)^{3/2}\left(m\wedge n\right)},$
@@
where we have used Lemma B.8 to bound $\sigma_{1}$ and $\sigma_{r}$, and $T\gtrsim L_{\phi}\log(d/\delta)$ (see the end of the proof of Lemma D.3) in the last inequalities. Since $\frac{4L^{2}}{\|M\|_{\max}}\geq\frac{\left(\|M\|_{\max}+\sigma\right)^{2}}{\|M\|_{\max}}\geq\|M\|_{\max}+\sigma\geq\sigma$, the first error term dominates all the others, and we can take
$K_{0}:=C^{\prime}\frac{L^{2}}{\|M\|_{\max}}\frac{\mu^{6}\kappa^{4}r^{3}\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}$
@@
This would yield an error bound whose dominant term scales with $\|\psi_{\pi}\|_{\Lambda}^{-1}\sqrt{\frac{\log\left(1/\delta\right)}{T}}$ and would echo the linear bandit upper bound of Theorem C.7, that we have demonstrated is instance-optimal. Although numerical experiments suggest this result holds, we were unsuccessful in our proof attempts. Notably, one could write $\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}^{2}=\|\psi_{\pi}\|_{\Lambda^{-1}}^{2}+(\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}^{2}-\|\phi_{\pi}\|_{\Lambda^{-1}}^{2})+(\|\phi_{\pi}\|_{\Lambda^{-1}}^{2}-\|\psi_{\pi}\|_{\Lambda^{-1}}^{2})$ and attempt to prove that the two rightmost terms converge to 00 with high probability. However, numerical experiments suggest that these terms do not both vanish in general, but rather cancel out, and thus a more careful analysis is required.
By leveraging Lemma D.12, we can however obtain the following instance-dependent bound, with a dependency on $\omega_{\min}$ instead of the entire behavior policy. Theorem 6.2 will follow from this result.
@@
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|$ $\leq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\sigma^{2}\log(16/\delta)}{\left(1-\alpha\right)T}}+K_{0}\left(\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}+\frac{\log^{5/4}\left(\left(m+n\right)T/\delta\right)}{T^{3/4}}\right),$
@@
$K_{0}\lesssim\frac{L^{2}}{\|M\|_{\max}}\frac{\mu^{6}\kappa^{4}r^{3}\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}.$
@@
$\|\phi_{\pi}\|^{2}_{\Lambda_{\phi}^{-1}}\leq\frac{\|\phi_{\pi}\|^{2}_{2}}{\omega_{\min}}\leq\frac{1}{\omega_{\min}}\left(\|\psi_{\pi}\|^{2}_{2}+3\epsilon_{\textup{Sub-Rec}}\right).$
Under the event $A_{\delta}$ defined in the proof of Lemma D.5, we consequently also have (using $\frac{1}{\sigma_{r}}\leq\frac{\mu^{2}\kappa r}{\sqrt{mn}\|M\|_{\max}}$)
$\|\psi_{\pi}\|^{2}_{\Lambda_{\phi}^{-1}}\leq\frac{1}{\omega_{\min}}\left(\|\psi_{\pi}\|^{2}_{2}+\frac{C}{\sqrt{T}}\frac{L}{\|M\|_{\max}}\mu^{3}\kappa^{2}r^{3/2}\sqrt{\frac{(m+n)}{\omega_{\min}mn\left(m\wedge n\right)}}\log^{3/2}\left(\frac{\left(m+n\right)T}{\delta}\right)\right)$ (31)
for some universal constant $C>0.$ Since $\sqrt{x+y}\leq\sqrt{x}+\sqrt{y}$ for every $x,y\geq 0$, this implies
$\|\phi_{\pi}\|_{\Lambda_{\phi}^{-1}}\leq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}+C^{1/2}\sqrt{\frac{L}{\|M\|_{\max}}}\mu^{3/2}\kappa r^{3/4}\omega_{\min}^{-3/4}\left(\frac{m+n}{mn\left(m\wedge n\right)}\right)^{1/4}\frac{\log^{3/4}\left(\left(m+n\right)T/\delta\right)}{T^{1/4}}.$ (32)
Combining this result with Proposition D.2151515Note that there is no need for a union bound as the high probability bound (32) is true under $A_{\delta}$. and noting that $\sigma\sqrt{\frac{L}{\|M\|_{\max}}}\leq\frac{\sigma L}{\|M\|_{\max}}\leq\frac{L^{2}}{\|M\|_{\max}}:=A$, we get that with probability larger than $1-\delta,$
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|$ $\leq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\sigma^{2}\log(16/\delta)}{\left(1-\alpha\right)T}}+C_{1}A\sqrt{\log(e/\delta)}\mu^{3/2}\kappa r^{3/4}\omega_{\min}^{-3/4}\left(\frac{m+n}{mn\left(m\wedge n\right)}\right)^{1/4}\frac{\log^{3/4}\left(\left(m+n\right)T/\delta\right)}{T^{3/4}}$
$+C_{2}A\mu^{6}\kappa^{4}r^{3}\frac{\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}$
for some universal constants $C_{1},C_{2}>0$. Since $\sqrt{\log(e/\delta)}\mu^{3/2}\kappa r^{3/4}\omega_{\min}^{-3/4}\left(\frac{m+n}{mn\left(m\wedge n\right)}\right)^{1/4}\log^{3/4}\left(\left(m+n\right)T/\delta\right)\lesssim\mu^{6}\kappa^{4}r^{3}\frac{\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\log^{5/4}\left(\left(m+n\right)T/\delta\right)$, this concludes the proof.
@@
When $M$ and $\omega$ are homogeneous, $K_{0}\lesssim\frac{(\|M\|_{\max}\vee\sigma)^{2}}{\|M\|_{\max}}mn$, and if we additionally have $\|M\|_{\max}=\Theta(\sigma),$ $K_{0}\lesssim\sigma mn.$
When both the context distribution and the behavior policy are uniform, the upper bound of Theorem D.7 corresponds to the conjectured instance-dependent upper bound mentioned in Remark D.6. Indeed, Lemma D.11 ensures that in this case, $\Lambda=\frac{1}{mn}I_{d},$ and thus $\|\psi_{\pi}\|_{\Lambda^{-1}}=\sqrt{mn}\|\psi_{\pi}\|_{2}=\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}.$
We now explain how to reformulate the statement of Theorem D.7 as a sample complexity result and retrieve the statement of Theorem 6.2.
@@
$\left|v^{\pi}-\hat{v}_{\text{\tt RS-PE}}\right|\leq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\sigma^{2}\log(16/\delta)}{\left(1-\alpha\right)T}}+K_{0}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T}+K_{0}\frac{\log^{5/4}\left(\left(m+n\right)T/\delta\right)}{T^{3/4}}$
@@
$\frac{\varepsilon}{3}\geq\frac{\|\psi_{\pi}\|_{2}}{\sqrt{\omega_{\min}}}\sqrt{\frac{2\sigma^{2}\log(16/\delta)}{\left(1-\alpha\right)T}}\text{,}\quad\frac{\varepsilon}{3}\geq K_{0}\frac{\log^{3}\left(\left(m+n\right)T/\delta\right)}{T},\quad\text{and}\quad\frac{\varepsilon}{3}\geq K_{0}\frac{\log^{5/4}\left(\left(m+n\right)T/\delta\right)}{T^{3/4}}.$
@@
$T\gtrsim\frac{\sigma^{2}\|\psi_{\pi}\|^{2}_{2}}{\omega_{\min}\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{K_{0}^{4/3}}{\varepsilon^{4/3}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)\!$ (33)
is enough to satisfy all three inequalities, where we recall that $K_{0}$ has been defined in (28).
Furthermore, the conditions on the number of samples of Theorem D.7, namely $T\gtrsim\frac{L^{2}\kappa^{2}(m+n)}{\sigma_{r}^{2}\omega_{\min}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$ and $T\gtrsim\frac{1}{\omega_{\min}}\log\left(16d/\delta\right)$, are both implied by (33) when $\varepsilon\leq\|M\|_{\max}$ because this choice implies $T\gtrsim\frac{K_{0}}{\|M\|_{\max}}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$ (to see that the second condition is verified, note that $\frac{K_{0}}{\|M\|_{\max}}\gtrsim\frac{(m+n)^{2}}{\omega_{\min}\sqrt{mn}\left(m\wedge n\right)}\gtrsim\frac{1}{\omega_{\min}}$).
@@
$K_{1}:=K_{0}^{4/3}\lesssim\left(\frac{L^{2}}{\|M\|_{\max}}\frac{\mu^{6}\kappa^{4}r^{3}\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}\right)^{4/3}.$ (34)
We finally note that by Remark D.8, when $M$ and $\omega$ are homogeneous and $\|M\|_{\max}=\Theta(\sigma)$, $K_{1}\lesssim\left(\sigma mn\right)^{4/3}.$
Minimax sample complexity guarantee. Let us note that by (40), $\|\psi_{\pi}\|^{2}_{2}\leq\mu^{2}r\frac{m+n-\mu^{2}r}{mn}$, which implies that $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt RS-PE}}|\leq\varepsilon$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n-\mu^{2}r\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{K_{1}}{\varepsilon^{4/3}}\log^{3}\left(\frac{(m+n)T}{\delta}\right).$
By adapting the previous analysis with a more refined argument, the dependency in $\varepsilon$ and the model parameters of the second term can be improved. This is what we demonstrate now.
@@
Let $\varepsilon,\delta\in(0,1)$. With the choices $T_{1}=\lfloor\alpha T\rfloor$ and $\tau\leq r(m\wedge n)^{-1}(\omega_{\min}/\omega_{\max})\log(16d/\delta)$, $|v^{\pi}-\hat{v}^{\pi}_{\text{\tt RS-PE}}|\leq\varepsilon$ with probability larger than $1-\delta$ as soon as
$T\gtrsim\frac{\sigma^{2}\mu^{2}r\left(m+n\right)}{\omega_{\min}mn\,\varepsilon^{2}}\log\left(\frac{e}{\delta}\right)+\frac{K_{0}}{\varepsilon}\log^{3}\left(\frac{(m+n)T}{\delta}\right)$
where
$K_{0}\lesssim\frac{L^{2}}{\|M\|_{\max}}\frac{\mu^{6}\kappa^{4}r^{3}\left(m+n\right)^{2}\sqrt{mn}}{\left(\omega_{\min}mn\right)^{2}\left(m\wedge n\right)}.$ (35)
@@
When $M$ and $\omega$ are homogeneous and $\|M\|_{\max}=\Theta(\sigma)$, $K_{0}\lesssim\sigma mn.$
Proof of Theorem D.9.
Combining (31) with the bound on $\|\psi_{\pi}\|^{2}_{2}$ entails
$\|\psi_{\pi}\|^{2}_{\Lambda_{\phi}^{-1}}$ $\leq\frac{1}{\omega_{\min}}\left(\frac{\mu^{2}r\left(m+n\right)}{mn}+\frac{C}{\sqrt{T}}\frac{L}{\|M\|_{\max}}\mu^{3}\kappa^{2}r^{3/2}\sqrt{\frac{m+n}{mn\left(m\wedge n\right)\omega_{\min}}}\log^{3/2}\left(\frac{\left(m+n\right)T}{\delta}\right)\right)$
$=\frac{\mu^{2}r\left(m+n\right)}{\omega_{\min}mn}\left(1+\frac{C}{\sqrt{T}}\frac{L}{\|M\|_{\max}}\mu\kappa^{2}r^{1/2}\sqrt{\frac{mn}{\left(m+n\right)\left(m\wedge n\right)\omega_{\min}}}\log^{3/2}\left(\frac{\left(m+n\right)T}{\delta}\right)\right)$ (36)